OpenStudy (anonymous):

the correct simplification of the expression -5x2(4x - 6x2 - 3).

OpenStudy (igreen):

Distribute it into the parenthesis. \(-5x^2 \times 4x\) \(-5x^2 \times -6x^2\) \(-5x^2 \times -3\)

OpenStudy (igreen):

Can you multiply those together?

OpenStudy (anonymous):

yes hold on

OpenStudy (igreen):

Remember, when we distribute we distribute the number(s) outside of the parenthesis to all the numbers inside the parenthesis.

OpenStudy (anonymous):

so would it be -30x^4+20x^3-15x^2?

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

That wouldn't be -30..-5 * -6 isn't -30.

OpenStudy (igreen):

Also, it wouldn't be 20 or -15.

OpenStudy (igreen):

Everything else is correct, just the whole numbers. -5 * 4 = ? -5 * -6 = ? -5 * -3 = ?

OpenStudy (anonymous):

then its -11x^4-x^3-8x^2

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

. . . No.. What's -5 * 4?

OpenStudy (igreen):

Only thing wrong here is your signs: \(-30x^4+20x^3-15x^2\) It should be: \(30x^4 - 20x^3 + 15x^2\)

OpenStudy (anonymous):

ohh i gotchuthank you